OpenStudy (vishweshshrimali5):

@rayford here is your question simplify \[\cfrac {2x^2-50}{x-5}\]

OpenStudy (vishweshshrimali5):

take out 2 common from the numerator \[\cfrac {(2)(x^2-25)}{x-5}\] you know that \[\color{red}{a^2 - b^2 = (a+b)(a-b)}\] now using this identity \[\large x^2 - 25 = \color {red} {(x+5)} \color {green}{(x-5)}\] Now you solve the fraction further @rayford